KjSaif (kjsaif):

Which relationship between x and y in the equation shows a proportional relationship?

KjSaif (kjsaif):

x+2=y+4

KjSaif (kjsaif):

y=4x+2

OpenStudy (phi):

a line that goes through (0,0)

KjSaif (kjsaif):

Which is the origin

KjSaif (kjsaif):

y=8x

OpenStudy (phi):

can you figure it out ?

KjSaif (kjsaif):

\[y=\frac{ x+4 }{ 2 }\]

KjSaif (kjsaif):

I'll try to figure it out

OpenStudy (phi):

replace x with 0 in each equation. only one will "give" y=0

KjSaif (kjsaif):

Is the answer y=8x?

KjSaif (kjsaif):

I believe it's y=8x

OpenStudy (phi):

yes, because y= 8x with x=0 gives y= 8*0 and y= 0

KjSaif (kjsaif):

Thanks @phi!

KjSaif (kjsaif):

good luck too 100ss! @phi